OpenStudy (anonymous):

determine the antiderivative of (x^2) sqrt (25-x^2) dx show steps plz

sam (.sam.):

\[\int\limits x^2 \sqrt{25-x^2}dx\]?

sam (.sam.):

Trig sub is what you need

sam (.sam.):

x=5sin(u) dx=5cos(u)du

sam (.sam.):

x/5=sinu u=sin^(-1)x/5

sam (.sam.):

\[25\int\limits 25 \sin ^2(u) \cos ^2(u) \, du\]

sam (.sam.):

oh man long step ahead

OpenStudy (anonymous):

@Sam I am trying to teach myself integration. How could you tell you needed a trig substitution ?